# Direct Data-Driven Predictive Control for a Three-dimensional Cable-Driven Soft Robotic Arm

### v1

@@ IV-A Setup and data collection
We validated the proposed method on a self-developed cable-driven soft robotic arm. The experimental setup is shown in Fig. 3. The arm is actuated by three cables, with a total effective length of 90 mm. End-effector positions were tracked in real time using a 12-camera OptiTrack FLEX13 motion capture system. Data collection and online control were executed on a host PC, with the DeePC framework implemented in MATLAB. Optimal control commands were transmitted via serial communication to a low-level controller, which drove the stepper motors to adjust cable lengths.
To construct a nonparametric representation, input–output data were collected. Here, $u^{d}$ denotes the incremental length of the three actuation cables, while $y_{d}=(p_{x},p_{y},p_{z})$ represents the end-effector position in task space. During data acquisition, ramp-and-hold excitations were applied to the low-level controller, and the corresponding end-effector positions were continuously recorded. The resulting dataset was then arranged into a Hankel matrix as described in Section III. When sufficiently rich, this data supports an accurate prediction of system behavior.
@@ IV-B Fixed-point reaching control
To evaluate DeePC’s ability to accurately drive the soft arm to specified target positions, we conducted a multi-stage reference tracking experiment. The bending angle $\phi_{b}$ and bending direction $\gamma_{g}$ were commanded to follow stepwise reference trajectories across three stages: (I) $\phi_{b}=20^{\circ},\ \gamma_{g}=0^{\circ}$; (II) $\phi_{b}=40^{\circ},\ \gamma_{g}=60^{\circ}$; (III) $\phi_{b}=60^{\circ},\ \gamma_{g}=120^{\circ}$. Here, $\phi_{b}$ and $\gamma_{g}$ are computed from the tip position using the constant curvature model [28].
@@ IV-C Trajectory tracking control
In this experiment, the arm was tasked to track a circular trajectory in task space. The reference trajectory was defined as a circle of radius $r$, discretized into a sequence of waypoints. For the baseline, control inputs were computed directly from the geometric model. For DeePC, the control inputs were constrained to $u\in[0,90]$, with $T_{\mathrm{ini}}=20$ and $N=30$. The weighting matrices were set to $Q=10\cdot I_{2\times 2}$ and $R=2\times 10^{-3}$, while the regularization parameters were chosen as $\lambda_{g}=300$ and $\lambda_{y}=1,000$.
For comparison, we implemented a baseline model-based controller derived from the constant-curvature assumption [25]. In this framework, the relationship between the backbone curvature of the arm and the cable length is expressed through idealized geometric relations. Specifically, for the $i$-th cable in a section:
@@
As shown in Fig. 5, DeePC achieved precise tracking with minimal deviation from the reference trajectory, both in 3D space (Fig. 5(a)) and in the xxyy projection (Fig. 5(b)). The DeePC trajectory closely overlapped the reference circle, maintaining a consistent radius and smooth curvature. In contrast, the baseline approach produced a distorted path: in 3D, the baseline trajectory sagged below the desired plane, while in the xxyy projection the circle was compressed into an irregular, triangular-like shape. These distortions reflect the limitations of constant-curvature geometric assumptions, which cannot capture the soft arm’s nonlinear bending and torsional behaviors.
@@
The physical bending configurations of the soft arm during circular trajectory tracking are illustrated in Fig. 6. The sequential snapshots show that DeePC successfully generated smooth, coordinated cable motions, enabling the arm to trace the desired circular path in real time.

### v2

@@ IV-A Setup and data collection
We validated the proposed method on a self-developed cable-driven soft robotic arm. The experimental setup is shown in Fig. 3. The arm is actuated by three cables, with a total effective length of 90 mm. End-effector positions were tracked in real time using a 12-camera OptiTrack FLEX13 motion capture system. Data collection and control were executed on a host PC, with the DeePC framework implemented in MATLAB. Optimal control commands were transmitted via serial communication to a low-level controller, which drove the stepper motors to adjust cable lengths.
To construct a nonparametric representation, input–output data were collected. Here, $u^{\mathrm{d}}$ denotes the incremental length of the three actuation cables, while $y^{\mathrm{d}}=(p_{x},p_{y},p_{z})$ represents the end-effector position in task space. During data acquisition, ramp-and-hold excitations were applied to the low-level controller, and the corresponding end-effector positions were continuously recorded. The resulting dataset was then arranged into a Hankel matrix as described in Section III. When sufficiently rich, this data supports an accurate prediction of system behavior.
@@ IV-B Fixed-point reaching control
To evaluate DeePC’s ability to accurately drive the soft arm to specified target positions, we conducted a multi-stage reference tracking experiment. The bending angle $\phi_{b}$ and orientation angle $\gamma_{g}$ were commanded to follow stepwise reference trajectories across three stages: (I) $\phi_{b}=20^{\circ},\ \gamma_{g}=0^{\circ}$; (II) $\phi_{b}=40^{\circ},\ \gamma_{g}=60^{\circ}$; (III) $\phi_{b}=60^{\circ},\ \gamma_{g}=120^{\circ}$. Here, $\phi_{b}$ and $\gamma_{g}$ are computed from the tip position using the constant curvature model [12].
@@ IV-C Trajectory tracking control
In this experiment, the arm was tasked to track a circular trajectory in task space. The reference trajectory was defined as a circle of radius $r$, discretized into a sequence of waypoints. For the baseline, control inputs were computed directly from the geometric model. For DeePC, the control inputs were constrained to $u\in[0,90]$, with $T_{\mathrm{ini}}=20$ and $N=30$. The weighting matrices were set to $Q=10\cdot I_{3\times 3}$ and $R=2\times 10^{-3}\cdot I_{3\times 3}$, while the regularization parameters were chosen as $\lambda_{g}=300$ and $\lambda_{y}=1,000$.
For comparison, we implemented a baseline model-based controller derived from the constant-curvature assumption [20]. In this framework, the relationship between the backbone curvature of the arm and the cable length is expressed through idealized geometric relations. Specifically, let $\kappa_{c,i}$ and $R_{c,i}$ denote the curvature and radius of curvature of the $i$-th cable, respectively, and let $l_{i}$ denote its length. For the $i$-th cable in a section:
@@
As shown in Fig. 5, DeePC achieved precise tracking with minimal deviation from the reference trajectory, both in 3D space (Fig. 5(a)) and in the xxyy projection (Fig. 5(b)). The DeePC trajectory closely overlapped the reference circle, maintaining a consistent radius and smooth curvature. In contrast, the baseline approach produced a distorted path: in 3D, the baseline trajectory sagged below the desired plane, while in the xxyy projection the circle was compressed into an irregular, triangular-like shape. These distortions reflect the limitations of constant-curvature geometric assumptions, which cannot capture the soft arm’s nonlinear bending and torsional behaviors. The physical bending configurations of the soft arm during circular trajectory tracking are illustrated in Fig. 6. The sequential snapshots show that DeePC successfully generated smooth, coordinated cable motions, enabling the arm to trace the desired circular path in real time.